OpenStudy (anonymous):

Please help.. Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1)

OpenStudy (anonymous):

from the first equation obtain the value of y y = 3 - 2x substitute this value in the second equation -3x + 4(3-2x) = 1 now solve for x and then find y by plugging the value of x into y = 3 - 2x

OpenStudy (anonymous):

Can you take it from here?

OpenStudy (texaschic101):

you understand what akitav is saying ?